OpenStudy (anonymous):

Complete the table of values for the following equation. 8x + 6y = 26

OpenStudy (anonymous):



OpenStudy (anonymous):

help i dont understand :,(

OpenStudy (anonymous):

@andycap You have there the equation and the values of x and y on the table. If x is missing, all you need to do is to replace y that is given on the table. I'll do the first one for you. :) \[8x+6y=26\] y is given so replace it on the solution. \[8x+42=26\] Transpose 42. \[8x=26-42\] \[8x=-16\] Divide both sides by 8. \[x=-2\] Try those on the other examples then I'll check. :)

OpenStudy (anonymous):

@andycap Did you get it? :)

OpenStudy (anonymous):

to tell you the thruth no,,, im not that good on math sorry if I didnt get it :S,,, And the problem is that i have a bunch of those questions

OpenStudy (anonymous):

im trying to understand haha

OpenStudy (anonymous):

@andycap Alright. I'll help you with the others. \[8(4)+6y=26\]\[32+6y=26\]\[6y=26-32\]\[6y=-6\]\[y=-1\] For the last one, \[8(10)+6y=26\]\[80+6y=26\]\[6y=26-80\]\[6y=-54\]\[y=-9\] Those are the answers. :)

OpenStudy (anonymous):

@andycap Did you get it now? :)

OpenStudy (anonymous):